OpenStudy (anonymous):

The value(in cents) of the change in a purse that contains twice as many nickles as pennies, four more dimes than nickels, and as many quarters as dimes and nickels combined. p=number of pennies express the given quantity in terms of p

OpenStudy (mathstudent55):

number of pennies = p value of p pennies (in cents) = (p)(1) = p "twice as many nickles as pennies" number of nickles = 2p value of 2p nickels = (2p)(5) = 10p "four more dimes than nickels" number of dimes = 2p + 4 value of 2p + 4 dimes (in cents) = (2p + 4)(10) = 20p + 40 "as many quarters as dimes and nickels combined" number of quarters = (2p + 4) + (2p) = 4p + 4 value of 4p + 4 quarters (in cents) = (4p + 4)(25) = 100p + 100 Total value = (p) + (10p) + (20p + 40) + (100p + 100) Total value = (131p + 140) cents

OpenStudy (anonymous):

wow thank you

OpenStudy (mathstudent55):

Do it one step at a time. It's easier than you think. You're welcome.